# ELIP: Enhanced Visual-Language Foundation Models for Image Retrieval

### v1

@@ Abstract
The objective in this paper is to improve the performance of text-to-image retrieval. To this end, we introduce a new framework that can boost the performance of large-scale pre-trained vision-language models, so that they can be used for text-to-image re-ranking. The approach, Enhanced Language-Image Pre-training (ELIP), uses the text query to predict a set of visual prompts to condition the ViT image encoding. ELIP can easily be applied to the commonly used CLIP/SigLIP and the state-of-the-art BLIP-2 architectures. To train the architecture with limited computing resources, we develop a ‘student friendly’ best practice involving global hard sample mining, and selection and curation of a large-scale dataset. On the evaluation side, we set up two new out-of-distribution benchmarks, Occluded COCO and ImageNet-R, to assess the zero-shot generalisation of the models to different domains. Benefiting from the novel architecture and data curation, experiments show our enhanced network significantly boosts CLIP/SigLIP performance and outperforms the state-of-the-art BLIP-2 model on text-to-image retrieval.
Figure 1: The ELIP architecture. Left: We propose a novel architecture that can be applied to pre-trained and frozen vision-language foundation models, such as CLIP, SigLIP and BLIP-2, to enhance their text-to-image retrieval performance. The key idea is to use the text query to define a set of visual prompt vectors that are incorporated into the image encoder to make it aware of the query when generating the embedding. An MLP maps from the text space to the visual space of the input to the ViT encoder. The architecture is lightweight, and our data curation strategies enables efficient and effective training with limited resources. Right: In this retrieval example from the COCO benchmark, the top-$k$ ($k$=100) images are re-ranked by our ELIP model for the text query: ‘People on bicycles ride down a busy street’. The ground truth image matching the query is not in the top-5 ranked images in the initial CLIP ranking, but is ranked top-1 (highlighted in the dashed box) by the re-ranking.
@@ 1 Introduction
In this paper, we consider the classic problem of text-to-image retrieval, that aims to rank image instances based on their relevance to a text query. Effective retrieval generally includes two stages: the first stage provides an initial ranking in a fast and efficient manner, while the second stage—referred to as re-ranking–refines this ranking by re-computing the relevance scores between the text query and each of the top-ranked candidates with a more expensive re-ranking model.
@@
In this paper, our primary contribution focuses on the second stage of the retrieval pipeline, the re-ranking. Specifically, our goal is to enhance the performance of off-the-shelf vision-language foundation models, so that they can be re-purposed for re-ranking the top-$k$ candidates from the fast retrieval process. The approach we develop, termed Enhanced Language-Image Pre-training (ELIP), requires only a few trainable parameters, and the training can be conducted efficiently with ‘student friendly’ resources and data. We demonstrate that ELIP can boost the performance of the pre-trained CLIP [65], SigLIP [93], and BLIP-2 [47] image-text models.
@@
To achieve this goal, first, we introduce a lightweight, text-guided visual prompting module. In our approach, as illustrated in Figure 1, a query text is mapped to a set of visual prompt vectors [41], which is then concatenated with the [CLS] and patch embeddings of the image encoder. These augmented embeddings are then passed into the frozen vision encoder to re-compute the image representation. The resulting image embedding is aware of the text conditioning and this enhances its performance in re-ranking.
@@
As a second contribution, we address the problem of training the model. There are two major challenges in training large vision-language models: first, data size – to enable a strong generalisation capability it is necessary to train on millions or billions of images, but this is expensive; second, the batch size – to enhance the model’s discriminative capability it is important to train at a large batch size, but that requires a large number of GPUs. We develop here a best practice by introducing strategies to select and curate a training dataset with maximum information, and group hard samples together in a batch to make training with small batch size effective.
@@
To evaluate the re-ranking performance of the ELIP models, we use the standard COCO [50] and Flickr30k [63] text-to-image retrieval benchmarks. As a further challenge, we also evaluate the generalisation of the ELIP boosted models on out-of-distribution domains. To do so, we repurpose the Occluded COCO [44] and ImageNet-R [33] datasets to be used for text-to-image retrieval benchmarks.
@@
In summary, we have made the following four contributions: First, we propose a novel architecture to improve text-based image retrieval on large pre-trained vision-language models, including the most popular CLIP/SigLIP architectures and the state-of-the-art BLIP-2 architecture. Second, we propose a best practice for training our architecture efficiently with limited resources. Third, to evaluate the generalisation capability of text-to-image retrieval models to different out-of-distribution domains, we set up two new benchmarks of text-to-image retrieval, Occluded COCO and ImageNet-R. Fourth, we demonstrate that ELIP significantly improves CLIP/SigLIP, and outperforms the state-of-the-art BLIP-2 model, on both standard benchmarks, COCO and Flickr, and on the new OOD benchmarks.
@@ 2 Related Work
Text-to-Image Retrieval is a fundamental and much researched task in cross-modal learning [45, 12, 39, 25, 96, 65, 46, 15, 81, 91, 16, 11, 13, 14, 19, 20, 22, 28, 35, 38, 42, 43, 48, 52, 53, 72, 75, 79, 80, 82, 83, 84, 85, 89, 92, 95, 97, 98, 99]. However, large-scale pre-trained vision language models such as CLIP [65, 36], SigLIP [93], ALIGN [40], and BLIP-2 [47], that have powerful zero-shot capabilities have now become the de facto method for open-set text based image retrieval, The most recent work [71] gives a slight improvement over BLIP-2 by incorporating the output of an object detector or annotations of detection bounding boxes. This succeeds in overcoming the failure cases where small but semantically important objects in an image are not properly understood by the model. We compare to this model, and exceed its performance.
@@
Post-Retrieval Re-ranking. For single modality image retrieval, where the query is an image, there have been a series of works that have re-ranked the top-$k$ images from an initial ranking via classical computer vision algorithms, such as ‘query expansion’, ‘geometric verification’, or a combination of the two [37, 62, 18, 17, 77, 3], as well as via learning based algorithms [8, 32, 21, 74, 4]. Re-ranking algorithms have been relatively less explored in text-to-image retrieval [90, 64, 54]. [57] introduced a method for computing the similarity score between an image and a text query by estimating the log-likelihood of the text conditioned on the image. While this approach has demonstrated strong performance, it remains computationally expensive both during training and inference, making it a slow process. Our paper also focuses on the re-ranking stage – developing a more powerful version of visual-language foundation models to give a better ranking of images that are hard to distinguish by the original retrieval model.
@@
Multi-Modal Datasets. To obtain multi-modal foundation models with a strong capability of generalisation, it is important to train them on large-scale multi-modal datasets. Therefore, in recent years, there has been a significant increase in the number and scale of multimodal vision-language datasets that provide image-text pairs, such as COCO [50], SBU [61], Conceptual Captions [69], LAION [67], DataComp [27]. The increase in the size of multi-modal datasets enables training of more powerful visual-language foundation models. More recently, DataCompDR [78] utilises prior knowledge from large-scale pre-trained image captioning models to generate synthetic captions for DataComp images which results in less noisy captions than the datasets collected from the web, such as the original DataComp dataset. In our paper, we have experimented training our model using Conceptual Captions [69] and DataCompDR [78].
@@
Multi-Modal Data Curation. It is essential to conduct data curation for multi-modal datasets as it enables more efficient and effective training, especially for situations where the resources are limited. There have been continuous efforts in data curation, such as offline example-level data pruning [27, 6, 9, 40, 26, 87, 34, 56], offline cluster-level data pruning [1, 2, 73, 7, 30], and online data curation with model-based scoring [23, 51, 55, 58]. The most recent work, JEST [24], utilises a pair of learner model and reference model to select batches of data that is learnable by the model but has not yet been learned. This inspired us to select the most efficient batches to train our architecture on BLIP-2. Another series of works related to us is hard negative mining, which has been both explored in classical metric learning [5, 31, 59, 70, 86, 88] and modern contrastive learning [66, 76].
Efficient Visual-Languge Pre-Training. To pre-train visual-language foundation models from scratch more efficiently, [29, 49] drop image tokens in the pre-training process to enable faster training, and training on batches with more samples.
@@ 3 Preliminaries
Re-Ranking in Image Retrieval. Given an input query, the goal of a retrieval system is to rank all instances in a dataset $\Omega=\{I_{1},\dots,I_{n}\}$, based on their relevance to the query. In the case of text to image retrieval, the query is specified by text ($T$), and the ideal outcome is a set ($\hat{\Omega}$), with the relevant images being ranked higher than those that are not. In general, an effective retrieval system proceeds in two stages: the first stage provides an initial ranking in a fast and efficient manner, while the second stage—referred to as re-ranking–refines this ranking by recomputing the relevance scores between the text query and each of the TopK ranked candidates with a more powerful (and usually more expensive) ranking model. The K is selected such that in general there is a high recall for all the relevant images. In this paper, our novelty focuses on the second stage, that aims to re-rank the TopK candidates from the Stage-I results.
Visual Prompt Tuning (VPT) [41] is a method of improving the ViT image encoder by inserting additional learnable prompts into the transformer layers. It enables efficient adaptation of ViT, requiring only the few parameters of the learnable prompts to be trained. VPT has two different variants – VPT-Shallow and VPT-Deep. VPT-Shallow only inserts the additional visual prompts into the first Transformer layer, whereas for VPT-Deep, prompts are introduced at every transformer layer’s input space. We insert our generated set of visual prompt vector into the first transformer layer of ViT, which is similar to VPT-Shallow.
@@ 4 The ELIP Architecture
In this section we describe the ELIP text-to-visual prompt mapping network, that can be efficiently applied to adapt the commonly used CLIP/SigLIP architectures as well as the state-of-the-art BLIP-2 architectures for re-ranking. We first introduce the architecture of the network in Section 4.1, and the training/inference strategy on CLIP/SigLIP and BLIP-2 in Sections 4.2 and 4.3 respectively. We refer to the network applied to CLIP as ELIP-C, applied to SigLIP as ELIP-S and applied to BLIP-2 as ELIP-B.
Figure 2: Architecture of ELIP-C/ELIP-S. At training time, a batch of text-image pairs is fed into the architecture. The text feature is mapped to the visual embedding space as a set of prompt vectors via the MLP mapping network, and then guides the encoding of image feature. We use orange for the [CLS] token, light blue for patch tokens, and dark blue for the generated visual tokens from text. The architecture is trained with InfoNCE loss (for ELIP-C) and Sigmoid loss (for ELIP-S) such that the text feature is aligned with the corresponding re-computed image feature.
@@ 4.1 Text-Guided MLP Mapping Network
In this work, we propose a mapping network that projects the embedding of the text query into a set of prompt vectors within the visual embedding space. This set of prompt vectors is then incorporated as additional tokens into the first layer of the Vision Transformer (ViT) image encoder, used to re-compute the visual embeddings:
@@
where $T$ denotes the query text, which is first encoded with a pre-trained, frozen text encoder ($\Phi_{\text{t}}(\cdot)$) into $m+1$ embeddings. The [CLS] token is further fed into a trainable mapping network to generate the prompt vectors, which are concatenated with the $n+1$ image embeddings ($[x_{p}^{1},\dots,x_{p}^{n},x_{\texttt{CLS}}]$), and passed into the pre-trained, frozen visual encoder ($\Phi_{\text{v}}(\cdot)$). The architecture is shown in Figure 2 and Figure 3.
4.2 Training and Inference with ELIP-C/ELIP-S
Figure 2 shows the application of our architecture on CLIP/SigLIP.
Text-Guided Contrastive Training. At training time, we compute the dot product between the [CLS] token embedding of the text query ($t_{\texttt{CLS}}$), and the re-computed image features guided by the query text, i.e., $\{v_{1},\dots,v_{b}\}$ ($b$ denotes the batch size). For ELIP-C, we take a softmax over all the dot products, and train with the standard InfoNCE loss on the batches; For ELIP-S, we train with pairwise Sigmoid loss for each dot product. In Section 5.1, we provide more details on the batch selection scheme via global hard sample mining.
Re-Ranking at Inference Time. At inference time, for each text query, we first compute the similarity scores between the visual-language embedding, computed by the original CLIP model, to obtain an initial ranking of all images. We then select the top-$k$ candidates for further re-ranking, where, the visual features are re-computed by incorporating the prompted vectors from the mapping network. The final ranking is obtained via the dot product of the re-computed image features and the text feature.
Figure 3: Architecture of ELIP-B. Similar to the architecture on CLIP/SigLIP, the MLP Mapping Network maps the text feature to the visual embedding space. The only difference is the text guided image features are further fed into the Q-Former to cross-attend to the input text and then passed through the Image-Text Matching (ITM) Head to predict whether the image and text match or not. As the input image features to the ITM head have been changed, we also fine-tune the ITM head, which is also a lightweight MLP network. The network is fed pairs of text and positive/negative image features at training time, and is trained with binary cross entropy loss.

### v2

@@ Abstract
The objective in this paper is to improve the performance of text-to-image retrieval. To this end, we introduce a new framework that can boost the performance of large-scale pre-trained vision-language models, so that they can be used for text-to-image re-ranking. The approach, Enhanced Language-Image Pre-training (ELIP), uses the text query, via a simple MLP mapping network, to predict a set of visual prompts to condition the ViT image encoding. ELIP can easily be applied to the commonly used CLIP, SigLIP and BLIP-2 networks. To train the architecture with limited computing resources, we develop a ‘student friendly’ best practice, involving global hard sample mining, and curation of a large-scale dataset. On the evaluation side, we set up two new out-of-distribution (OOD) benchmarks, Occluded COCO and ImageNet-R, to assess the zero-shot generalisation of the models to different domains. The results demonstrate that ELIP significantly boosts CLIP/SigLIP/SigLIP-2 text-to-image retrieval performance and outperforms BLIP-2 on several benchmarks, as well as providing an easy means to adapt to OOD datasets.
Figure 1: The ELIP architecture. Left: We propose a novel architecture that can be applied to pre-trained and frozen vision-language foundation models, such as CLIP, SigLIP, SigLIP-2 and BLIP-2, to enhance their text-to-image retrieval performance. The key idea is to use the text query to define a set of visual prompt vectors that are incorporated into the image encoder to make it aware of the query when generating the embedding. An MLP maps from the text space to the visual space of the input to the ViT encoder. The architecture is lightweight, and our data curation strategies enable efficient and effective training with limited resources. Right: In this retrieval example from the COCO benchmark, the top-$k$ ($k$=100) images are re-ranked by our ELIP model for the text query: ‘People on bicycles ride down a busy street’. The ground truth image matching the query is not in the top-5 ranked images in the initial CLIP ranking, but is ranked top-1 (highlighted in the dashed box) by the re-ranking.
@@ 1 Introduction
This paper considers the problem of text-to-image retrieval, that aims to rank image instances based on their relevance to a text query. Effective retrieval generally includes two stages: the first stage provides an initial ranking in a fast and efficient manner, while the second stage—referred to as re-ranking–refines this ranking by re-computing the relevance scores between the text query and each of the top-ranked candidates with a more expensive model.
@@
Our primary contribution here focuses on the second stage of the retrieval pipeline, namely, the re-ranking. Specifically, our goal is to enhance the performance of off-the-shelf vision-language foundation models, so that they can be re-purposed for re-ranking the top-$k$ candidates from the fast retrieval process. The approach we develop, termed as Enhanced Language-Image Pre-training (ELIP), requires only a few trainable parameters, and the training can be conducted efficiently with ‘student friendly’ resources and data. We demonstrate that ELIP can boost the performance of the pre-trained CLIP [68], SigLIP [99], SigLIP-2 [81] and BLIP-2 [48] for cross-modal retrieval.
@@
To achieve this goal, we first introduce a lightweight, text-guided visual prompting module. As illustrated in Figure 1, a query text is mapped to a set of visual prompt vectors [42], that are then concatenated with the [CLS] and patch embeddings of the image encoder. These augmented embeddings are then passed into the frozen vision encoder to re-compute the image representation. The resulting image embedding is aware of the text conditioning and this enhances its performance in re-ranking.
@@
As a second contribution, we address the two major challenges in training large vision-language models: first, data size – to enable a strong generalisation capability it is necessary to train on millions or billions of images, but this is expensive; second, the batch size – to enhance the model’s discriminative capability it is important to train at a large batch size, but that requires a large number of GPUs. We develop here a best practice by introducing strategies to select and curate a training dataset with maximum information, and group hard samples together in a batch to make training with small batch size effective.
@@
To assess the re-ranking performance of our proposed ELIP models, we experiment on the standard COCO [50] and Flickr30k [66] text-to-image retrieval benchmarks. As a further challenge, we also evaluate the generalisation of the ELIP-boosted models on out-of-distribution domains. To do so, we repurpose the Occluded COCO [45] and ImageNet-R [34] datasets to be used for text-to-image retrieval benchmarks.
@@
In summary, we have made the following four contributions: First, we propose a novel architecture to improve text-based image retrieval on large pre-trained vision-language models, including the most popular CLIP/SigLIP architectures and the state-of-the-art BLIP-2 architecture. Second, we propose a best practice for training our architecture efficiently with limited resources. Third, to evaluate the generalisation capability of text-to-image retrieval models to different out-of-distribution domains, we set up two new benchmarks of text-to-image retrieval, Occluded COCO and ImageNet-R. Fourth, and most significantly, we demonstrate that ELIP substantially improves the image retrieval performance of CLIP and SigLIP architectures, and outperforms the state-of-the-art BLIP-2 architecture. Furthermore, it provides an efficient method to adapt these architectures to OOD datasets, again giving a tremendous boost with CLIP, SigLIP, SigLIP-2, and BLIP-2.
@@ 2 Related Work
Text-to-Image Retrieval is a fundamental and much researched task in cross-modal learning [46, 13, 40, 26, 102, 68, 47, 16, 86, 97, 17, 12, 14, 15, 20, 21, 23, 30, 36, 39, 43, 44, 49, 52, 53, 75, 78, 84, 85, 87, 88, 90, 91, 95, 98, 101, 103, 104, 105]. Large vision language models such as CLIP [68, 37], ALIGN [41], BLIP-2 [48], SigLIP [99] and SigLIP-2 [81] that have powerful zero-shot capabilities have now become the de facto method for open-set text-based image retrieval. The most recent work [74] gives a slight improvement over BLIP-2 by incorporating the output of an object detector or annotations of detection bounding boxes. This succeeds in overcoming the failure cases where small but semantically important objects in an image are not properly understood by the model. We compare to this model and show superior performance.
CIR and Universal Retrieval. In composed image retrieval (CIR) [55, 29, 5, 83], the query is specified by a composition of an image and text, with the text specifying how the image should be changed. For example, the query image may be of a dog lying down, and the query text may be ‘playing with a ball’. This composed query defines the target image to be retrieved from the gallery. This differs from our task, where the query is specified only by text, and the text alone defines the target image to be retrieved from the gallery. A more general setting is ‘universal retrieval’ [89, 56] where the query can be a combination of image, text, and instruction; and the target can be image alone, text alone, or image and text.
@@
Post-Retrieval Re-ranking. For single modality image retrieval, where the query is an image, there has been a series of works that have re-ranked the top-$k$ images from an initial ranking via classical computer vision algorithms, such as ‘query expansion’, ‘geometric verification’, or a combination of the two [38, 65, 19, 18, 80, 4], as well as via learning-based algorithms [10, 33, 22, 77, 6]. Re-ranking algorithms have been relatively less explored in text-to-image retrieval [96, 67, 57]. [60] introduced a method for computing the similarity score between an image and a text query by estimating the log-likelihood of the text conditioned on the image. While this approach has demonstrated strong performance, it remains computationally expensive both during training and inference, making it a slow process. Our paper also focuses on the re-ranking stage – developing a more powerful version of visual-language foundation models to give a better ranking of images that are hard to distinguish by the original retrieval model.
@@
Multi-Modal Datasets. To obtain multi-modal foundation models with a strong capability of generalisation, it is important to train them on large-scale multi-modal datasets. Therefore, in recent years, there has been a significant increase in the number and scale of multimodal vision-language datasets that provide image-text pairs, such as COCO [50], SBU [64], Conceptual Captions [72], LAION [70], DataComp [28]. The increase in the size of multi-modal datasets enables the training of more powerful visual-language foundation models. More recently, DataCompDR [82] utilises prior knowledge from large-scale pre-trained image captioning models to generate synthetic captions for DataComp images which results in less noisy captions than the datasets collected from the web, such as the original DataComp dataset. In our paper, we have experimented with training our model using Conceptual Captions [72] and DataCompDR [82].
@@
Multi-Modal Data Curation. It is essential to conduct data curation for multi-modal datasets as it enables more efficient and effective training, especially in situations where the resources are limited. There have been continuous efforts in data curation, such as offline example-level data pruning [28, 8, 11, 41, 27, 93, 35, 59], offline cluster-level data pruning [1, 2, 76, 9, 31], and online data curation with model-based scoring [24, 51, 58, 61]. The most recent work, JEST [25], utilises a pair of learner model and reference model to select batches of data that are learnable by the model but have not yet been learned. This inspired us to select the most efficient batches to train our architecture on BLIP-2. Another series of works related to us is hard negative mining, which has been both explored in classical metric learning [7, 32, 62, 73, 92, 94] and modern contrastive learning [69, 79].
@@ 3 Preliminaries
Re-Ranking in Image Retrieval. Given an input query, the goal of a retrieval system is to rank all instances in a dataset $\Omega=\{I_{1},\dots,I_{n}\}$, based on their relevance to the query. In the case of text-to-image retrieval, the query is specified by text ($T$), and the ideal outcome is a set ($\hat{\Omega}$), with the relevant images being ranked higher than those that are not. In general, an effective retrieval system proceeds in two stages: the first stage provides an initial ranking in a fast and efficient manner, while the second stage—referred to as re-ranking–refines this ranking by recomputing the relevance scores between the text query and each of the top-$k$ ranked candidates with a more powerful (and usually more expensive) ranking model. The $k$ is selected such that in general there is a high recall for all the relevant images. In this paper, our novelty lies on the second stage, that aims to re-rank the top-$k$ candidates from the first stage results.
Visual Prompt Tuning (VPT) [42] is a method of enhancing the ViT image encoder by inserting additional learnable prompts into the transformer layers. It enables efficient adaptation of ViT, requiring only the few parameters of the learnable prompts to be trained. VPT has two different variants – VPT-Shallow and VPT-Deep. VPT-Shallow only inserts the additional visual prompts into the first Transformer layer, whereas for VPT-Deep, prompts are introduced at every transformer layer’s input space. We insert our generated set of visual prompt vector into the first transformer layer of ViT, which is similar to VPT-Shallow.
@@ 4 The ELIP Architecture
In this section, we describe the ELIP text-to-visual prompt mapping network, that can be efficiently applied to adapt the commonly used CLIP/SigLIP architectures as well as the more sophisticated BLIP-2 architectures for re-ranking. We first introduce the architecture of the network in Section 4.1, and the training/inference strategy in Sections 4.2 and 4.3 respectively. We refer to the network applied to CLIP as ELIP-C, applied to SigLIP/SigLIP-2 as ELIP-S/ELIP-S-2, and applied to BLIP-2 as ELIP-B.
Figure 2: Architecture of ELIP-C / ELIP-S. At training time, a batch of text-image pairs is fed into the architecture. The text feature is mapped to the visual embedding space as a set of prompt vectors via the MLP mapping network and then guides the encoding of the image feature. We use color coding for the [CLS] token, patch tokens, and generated visual tokens from text. The architecture is trained with InfoNCE loss (for ELIP-C) and Sigmoid loss (for ELIP-S/ELIP-S-2), to align the text feature with the corresponding re-computed image feature.
@@ 4.1 Text-Guided MLP Mapping Network
Here, we propose a mapping network that projects the embedding of the text query into a set of prompt vectors within the visual embedding space. This set of prompt vectors is then incorporated as additional tokens into the first layer of the Vision Transformer (ViT) image encoder, used to re-compute the visual embeddings:
@@
where $T$ denotes the query text, which is first encoded with a pre-trained, frozen text encoder ($\Phi_{\text{t}}(\cdot)$) into $m+1$ embeddings. The [CLS] token is further fed into a trainable mapping network to generate the prompt vectors, which are concatenated with the $n+1$ image embeddings ($[x_{p}^{1},\dots,x_{p}^{n},x_{\texttt{CLS}}]$), and passed into the pre-trained, frozen visual encoder ($\Phi_{\text{v}}(\cdot)$). The MLP Mapping Network consists of 3 layers of linear layers with a GELU between every two linear layers. The ELIP architecture is shown in Figure 2 and Figure 3.
4.2 Training and Testing ELIP-C/ELIP-S
Text-Guided Contrastive Training. At training time, we compute the dot product between the [CLS] token embedding of the text query ($t_{\texttt{CLS}}$) and the re-computed image features guided by the query text, i.e., $\{v_{1},\dots,v_{b}\}$ ($b$ denotes the batch size). For ELIP-C, we train with the standard InfoNCE loss on the batches; For ELIP-S/ELIP-S-2, we train with pairwise Sigmoid loss. In Section 5.1, we provide more details on the batch selection scheme via global hard sample mining.
Re-Ranking at Inference Time. At inference time, for each text query, we first compute the similarity scores between the visual-language embedding, computed by the original CLIP/SigLIP model, to obtain an initial ranking of all images. We then select the top-$k$ candidates for further re-ranking, where, the visual features are re-computed by incorporating the prompted vectors from the mapping network. The final ranking is obtained via the dot product of the re-computed image features and the text feature.
Figure 3: Architecture of ELIP-B. Similar to the architecture on CLIP/SigLIP, the MLP Mapping Network maps the text feature to the visual embedding space. The only difference is the text-guided image features are further fed into the Q-Former to cross-attend the input text and then passed through the Image-Text Matching (ITM) Head to predict whether the image and text match or not. As the input image features to the ITM head have been changed, we also fine-tune the ITM head, which is a lightweight MLP network. The network is fed pairs of text and positive/negative image features at training time and is trained with binary cross entropy loss.